OpenStudy (anonymous):

ball 2 with a mass of 30 kg, is moving to the left at 10m/s. with what velocity should ball 1, with a mass of 10kg, move to the right and collide with 2, so that ball 1 rebounds with a velocity of 30m/s and ball 2 with velocity of 10m/s after the collision? assume the collision to be perfectly elastic

OpenStudy (anonymous):

You would use the momentum equation p=mv. In this situation, it is Mass(before) * Velocity(before) = M(after) * V(after). Ball 2(before) = 30*10 = 300. Ball 2(after) is the same, 10m/s and 30kg = 300. Ball 1(after) = 10kg * 30m/s = 300. Therefore, Ball 1(before) should also be 300.